Phyllis invested $ 12,000, a portion earning a simple interest rate of 4 1/2% per year and the rest earning a rate of 4% per year. After one year the total interest earned on these investment was $525. How much money did she invest at each rate??
Answer by checkley71(8403)

(x*.045)+(12000-x).04=525 or .045x+480-.04x=525 or .005x=525-480 or .005x=45 or
x=45/.005 or x=9000 invested @ 4.5% then 12000-9000=3000 invested @ 4%
proof 9000*.045+3000*.04=525 or 405+120=525 or 525=525